OpenStudy (anonymous):

Substitution help again anyone? Solve using substitution: -8x = -4 − y 3y -24x = 8 Select one: a. (1/4, -6) b. (-2, 12) c. Infinitely many solutions d. No solution

OpenStudy (anonymous):

\(-8x = -4 - y\) \(3y - 24x = 8\)

OpenStudy (anonymous):

Which way, you want to solve it?

OpenStudy (anonymous):

By using options you have, or you want to solve it step by step?

OpenStudy (anonymous):

options

OpenStudy (anonymous):

For example: I use \(b\) option : In that \(x = -2\) and \(y = 12\) given to you. Plug in x and y values in the equation given to you, and check whether Left Side equals Right Side or not.

OpenStudy (anonymous):

yes they are right?

OpenStudy (anonymous):

Right??

OpenStudy (anonymous):

\(-8(-2) = ??\)

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

\(-8(-2) \implies -8 \times -2 = ??\)

OpenStudy (anonymous):

24

OpenStudy (anonymous):

Are you serious?

OpenStudy (anonymous):

\(8 \times 2 = ??\)

OpenStudy (anonymous):

16 my bad haha

OpenStudy (anonymous):

\(-8 \times -2 = ??\)

OpenStudy (anonymous):

I dont know what it is if its not 16